OpenStudy (anonymous):

16 is what percent of 5

OpenStudy (anonymous):

x% of 5=16

OpenStudy (anonymous):

So, (x/100) *5=16

OpenStudy (anonymous):

16 is what percent of 5 = 16 / 5 = 3.2 Converting decimal to a percentage: 3.2 * 100 = 320%

OpenStudy (anonymous):

Now, slove for x